# An Analysis of Segment Anything 2

### v1

@@ Abstract
Video object segmentation (VOS) is a critical task in the development of video perception and understanding. The Segment-Anything Model 2 (SAM 2), released by Meta AI, is the current state-of-the-art architecture for end-to-end VOS. SAM 2 performs very well on both clean video data and augmented data, and completely intelligent video perception requires an understanding of how this architecture is capable of achieving such quality results. To better understand how each step within the SAM 2 architecture permits high-quality video segmentation, we pass a variety of complex video transformations through the architecture and measure the impact at each stage of the process. We observe that each progressive stage enables the filtering of complex transformation noise and the emphasis of the object of interest. Our contributions include the creation of complex transformation video datasets, an analysis of how each stage of the SAM 2 architecture interprets these transformations, and visualizations of segmented objects through each stage. By better understanding how each model structure impacts overall video understanding, VOS development can work to improve real-world applicability and performance tracking, localizing, and segmenting objects despite complex cluttered scenes and obscurations.
@@
1 Introduction
In determining the quality of VOS models, benchmarking and quantitative metrics can only tell so much. Although SAM 2 mask predictions are high quality using these metrics, they give us little information about how the model perceives objects. By observing how a model reacts to changes in input, we can determine how perception of VOS models like SAM 2 occurs. Each stage within the SAM 2 architecture alters object perception and, using a series of data transformations, it becomes evident where the model transitions from scene-oriented perception to object-oriented perception.
Understanding model perceptions can inform architecture decisions, loss decisions, and potential weaknesses. SAM 2 has flaws, including a weak tracking ability [9]. By probing SAM 2 with challenging data, we can identify potential problems and locate where improvements can be made. Identifying the layers of SAM 2 is vital to developing future improvements to video object segmentation models.
SAM 2’s modular design makes it easy to introduce observational positions throughout the model at the image encoding, cross-attention with memory for enriched image features, and cross-attention with sparse prompts for object attention refinement. Tracking embeddings during these observational positions as we introduce complex datasets designed to test the visual perception of models will enable insight into solving common challenges of designing VOS models for cluttered complex scenes and long-term obscurations.
Contributions
Our contributions are as follows.
@@ 1 Introduction
2 Related Works
2.1 Image Segmentation
@@ 2.2 Video Segmentation
A simplified architecture diagram for SAM 2 is shown in Figure 1. First, an input image is preprocessed and passed through an image encoder to find the image embeddings. These embeddings undergo two cross-attentional layers, first with the memory of previous mask/frame relations and second with prompt inputs, such as masks, bounding boxes, and point clicks. The mask predictions are decoded as a function of these cross-attentional layers. A 256-dimensional object pointer represents the current state of the object in relation to previous frames. The mask prediction and object pointer are encoded into memory and used for future predictions.
2.3 Datasets
@@
MOSE [4] is a popular video dataset similar to DAVIS that emphasizes object obscuration. In many samples, obscured objects will become partially or fully obscured in a section of the video to test the robustness of the VOS model. State-of-the-art models, particularly Cutie and SAM 2, have emphasized this dataset when developing their architectures and have great performance on these obscurations. Our datasets contain synthetic obscurations that are unreliable when the clean data contains natural obscurations. For this reason, we use DAVIS to create a dataset of videos with the necessary annotated long-term interjections.
3 Experimental Design
@@
3.1 Data Transformations
While the DAVIS video dataset is used as a baseline for all observed parameters, five complex transformations are applied to create new 500-video datasets. Stochastic resampling of DAVIS videos allows for random sample generation. We define the object of interest as the object which will be segmented, and obscuring object as a different object that is partially or fully overlapping the object of interest, and the context as the frame background that is not a part of either object.
3.1.1 Four-Frame Interjection
An object is selected from a video of interest, and a select number of frames from a separate, unrelated video are interjected in the middle of the video of interest [1], as shown in Figure 2. Thus, the object of interest is not present during the interjection period. SAM 2 is expected to do the following during the three stages of the interjection video:
@@ 3.1.4 Obscuration
3.1.5 Three-Object Overlay
@@ 3.2 Observational Positions
3.2.1 Position 1: Image Embeddings
@@
3.2.2 Position 2: Pixel/Memory Features
@@
3.2.3 Position 3: Pixel/Prompt Embeddings
@@
3.2.4 Position 4: Object Pointer
@@
3.2.5 Position 5: Memory Features
@@
4 Results
4.1 Obscuration Datasets
@@ 4.1.1 Obscuration
4.1.2 Three-Object Overlay
@@
4.2 Object Pointer Analysis
As described in the SAM 2 paper [8], the object pointer is a [1,256] vector used by the model to recall the position of the object in previous frames. It is created as an embedding of the object location and fed back during cross attention with memory. This raises the question of what exactly the object pointer represents and how it is perceived by the model. To determine if there is a spatial relationship between the object pointers and the object in the original frame, we explore basic models to decode the object position. A basic multilevel perceptron (MLP) model is trained on the entire DAVIS dataset with the [1,256] object pointer input and the ground truth [1,4] object bounding box output, representing [xmin, ymin, xmax, ymax]. This trained model is used to predict object location in the current frame using only the object pointer. Figure 14 shows the projection of the object pointer resulting in the same frames shown in Figures 7, 8, 11, and 12, respectively.
@@ 4.2 Object Pointer Analysis
4.3 Interjection Datasets
Interjection datasets refer to the datasets for which there is a four-frame interjection period (four-frame interjection, object removal, and context removal) as well as the clean DAVIS dataset as a baseline. It is important to note that the clean DAVIS and context-removal datasets maintain object presence throughout the entire video while the four-frame interjection and the object removal datasets do not have the object present for frames 12-15. Ideally, the explored features, defined in appendix A.1, should remain relatively constant for the frames that show the object. The features should increase significantly during the interjection period (frames 12-15) for the four-frame interjection and object removal datasets.
4.3.1 Position 1 - Image Embeddings
@@
In all three feature plots, the baseline clean DAVIS dataset is relatively constant as expected. Our metric applied to the four-frame interjection dataset (red line) increases significantly during the interjection period (frames 12-15), as each interjection frame is compared to the prefix frames (frames 0-11). On average, each interjection frame is further from the last known frame containing the object than two frames from the same video are from each other.
Our metrics applied to the context removal dataset (green line) follows an interesting pattern in the image embedding space. Ideally, these plots will be constant, like the clean DAVIS plots. Because the context is removed during the interjection period, the frames are very distant from earlier frames. The large jump in feature values in frame 12 reflect this frame dissimilarity. However, the object is still present. As such, the subsequent frames are being compared to the removed-context frames, resulting in a lower distance for the remainder of the interjection period. The short-term distance and ratio plots show an additional jump in frame 16 when the context returns.
Our metrics applied to the object removal dataset (orange line) experiences the inverse effect. These plots should match the four-frame interjection plots with a large jump during the interjection period. Because the background context of the frames in the interjection period remain, the frame is generally similar during this period. This is shown by the lack of a large increase during the interjection period. Instead. the feature values increase slightly, representing the small change in frame caused by the removed object.
4.3.2 Position 4 - Object Pointer
@@
The object pointer space is the ideal representation of the object separated from its surroundings. Analysis of the object pointer can create a nearly complete picture of the segmented object. Further analysis of the observed features and additional plots for all positions are shown in Appendix A.2
5 Conclusion
SAM 2, the current state-of-the-art video object segmentation model, employs a unique architecture that solely relies on cross-attention with memory. In order to improve real-world applicability of tracking, localization, and segmentation models, particularly on complex scenes, it is necessary to understand how each stage of the SAM 2 architecture interprets video. In this paper, we proposed five new datasets with complex transforms, each testing the robustness of SAM 2 on different challenges. We analyzed five different observational positions at each major stage within the model and determined at which stage the model is able to discern the object of interest from obscuring objects and the surrounding context. Primarily, the object pointer provides unique insight into where the model is spatially focused when writing to memory and when subsequent frames perform memory cross-attention. By visualizing how each architectural structure impacts the overall video understanding, VOS development can improve real-world applicability despite complex cluttered scenes and video transformations.
Acknowledgments
Appendix A Appendix
Additional results are included here, which contain complete visualizations and analysis of embedding locations across our complex transformed datasets.
A.1 Methods
@@
In real-world applications, it is useful to distinguish frames containing an object from frames in which the object is not present [1]. The computed metric can be used to create a classifier to separate visible frames from artificially interjected frames without the object. In this paper, we explore three features resulting from these computed regularized distances that may be used as a classifier. We aim to understand the nature of SAM 2 by quantifying the differences between clean and transformed videos in the equivalent observational positions.
A.1.1 Zeroth-Order L2 Comparison
@@
A.1.2 First-Order L2 Comparison
In certain cases, the most dissimilar frames from the same video may be less similar than two frames from different videos. As such, zeroth-order classification is imperfect. First-order classification looks at the ratio of the current zeroth-order distance to the previous distance. This allows the model to determine if an interjected frame is less similar to the original video than the original video is to itself, thus improving the classification accuracy. In this paper, we explore the short-term L2 ratio as a feature of each stage.
A.2 Extended Results
A.2.1 Position 1 - Image Embeddings
@@
A.2.2 Position 2 - Pixel/Memory Features
@@
A.2.3 Position 3 - Pixel/Prompt Embeddings
@@
A.2.4 Position 4 - Object Pointer
@@
A.2.5 Position 5 - Memory Features

### v2

# An Analysis of Data Transformation Effects on Segment Anything 2

@@ Abstract
Video object segmentation (VOS) is a critical task in the development of video perception and understanding. The Segment-Anything Model 2 (SAM 2), released by Meta AI, is the current state-of-the-art architecture for end-to-end VOS. SAM 2 performs very well on both clean video data and augmented data, and completely intelligent video perception requires an understanding of how this architecture is capable of achieving such quality results. To better understand how each step within the SAM 2 architecture permits high-quality video segmentation, a variety of complex video transformations are passed through the architecture, and the impact at each stage of the process is measured. It is observed that each progressive stage enables the filtering of complex transformation noise and the emphasis of the object of interest. Contributions include the creation of complex transformation video datasets, an analysis of how each stage of the SAM 2 architecture interprets these transformations, and visualizations of segmented objects through each stage. By better understanding how each model structure impacts overall video understanding, VOS development can work to improve real-world applicability and performance tracking, localizing, and segmenting objects despite complex cluttered scenes and obscurations.
Index Terms:
Artificial Intelligence, Computer Vision, Object Understanding, SAM 2, Video Data Transformation, Video Object Segmentation
@@
I Introduction
In determining the quality of VOS models, benchmarking and quantitative metrics can only tell so much. Although SAM 2 mask predictions are high quality using these metrics, they give little information about how the model perceives objects. By observing how a model reacts to changes in input, it can be determined how perception of VOS models like SAM 2 occurs. Each stage within the SAM 2 architecture alters object perception and, using a series of data transformations, it becomes evident where the model transitions from scene-oriented perception to object-oriented perception.
Understanding model perceptions can inform architecture decisions, loss decisions, and potential weaknesses. SAM 2 has flaws, including a weak tracking ability [1]. By probing SAM 2 with challenging data, potential problems can be identified and areas of improvement can be located. Identifying the layers of SAM 2 is vital to developing future improvements to video object segmentation models.
SAM 2’s modular design makes it easy to introduce observational positions throughout the model at the image encoding, cross-attention with memory for enriched image features, and cross-attention with sparse prompts for object attention refinement. Tracking embeddings during these observational positions with the introduction of complex datasets designed to test the visual perception of models will enable insight into solving common challenges of designing VOS models for cluttered complex scenes and long-term obscurations.
Contributions are as follows.
@@ I Introduction
II Related Works
II-A Image Segmentation
@@ II-B Video Segmentation
A simplified architecture diagram for SAM 2 is shown in Figure 1. First, an input image is preprocessed and passed through an image encoder to find the image embeddings. These embeddings undergo two cross-attentional layers, first with the memory of previous mask/frame relations and second with prompt inputs, such as masks, bounding boxes, and point clicks. Mask predictions are decoded as functions of these cross-attentional layers. A 256-dimensional object pointer represents the current state of the object in relation to previous frames. The mask prediction and object pointer are encoded into memory and used for future predictions.
II-C Datasets
@@
MOSE [7] is a popular video dataset similar to DAVIS that emphasizes object obscuration. In many samples, obscured objects will become partially or fully obscured in a section of the video to test the robustness of the VOS model. State-of-the-art models, particularly Cutie and SAM 2, have emphasized this dataset when developing their architectures and have great performance on these obscurations. The created datasets contain synthetic obscurations that are unreliable when clean data contain natural obscurations. For this reason, DAVIS is used to create a dataset of videos with the necessary annotated long-term interjections.
III Experimental Design
@@
III-A Data Transformations
While the DAVIS video dataset is used as a baseline for all observed parameters, five complex transformations are applied to create new 500-video datasets. Stochastic resampling of DAVIS videos allows for random sample generation. Object of interest is defined as the object which will be segmented, and obscuring object as a different object that is partially or fully overlapping the object of interest, and the context as the frame background that is not a part of either object.
III-A1 Four-Frame Interjection
An object is selected from a video of interest and a select number of frames from a separate, unrelated video are interjected in the middle of the video of interest [8], as shown in Figure 2. Thus, the object of interest is not present during the interjection period. SAM 2 is expected to do the following during the three stages of the interjection video:
@@ III-A4 Obscuration
III-A5 Three-Object Overlay
@@ III-B Observational Positions
III-B1 Position 1: Image Embeddings
@@
III-B2 Position 2: Pixel/Memory Features
@@
III-B3 Position 3: Pixel/Prompt Embeddings
@@
III-B4 Position 4: Object Pointer
@@
III-B5 Position 5: Memory Features
@@
IV Results
IV-A Obscuration Datasets
@@ IV-A1 Obscuration
IV-A2 Three-Object Overlay
@@
IV-B Object Pointer Analysis
As described in the SAM 2 paper [3], the object pointer is a [1,256] vector used by the model to recall the position of the object in previous frames. It is created as an embedding of the object location and fed back during cross attention with memory. This raises the question of what exactly the object pointer represents and how it is perceived by the model. To determine if there is a spatial relationship between the object pointers and the object in the original frame, basic models are explored to decode the object position. A basic multilevel perceptron (MLP) model is trained on the entire DAVIS dataset with the [1,256] object pointer input and the ground truth [1,4] object bounding box output, representing [xmin, ymin, xmax, ymax]. This trained model is used to predict object location in the current frame using only the object pointer. Figure 14 shows the projection of the object pointer resulting in the same frames shown in Figures 7, 8, 11, and 12, respectively.
@@ IV-B Object Pointer Analysis
IV-C Interjection Datasets
Interjection datasets refer to the datasets for which there is a four-frame interjection period (four-frame interjection, object removal, and context removal) as well as the clean DAVIS dataset as a baseline. It is important to note that the clean DAVIS and context-removal datasets maintain object presence throughout the entire video while the four-frame interjection and the object removal datasets do not have the object present for frames 12-15. Ideally, the explored features, defined in appendix A, should remain relatively constant for the frames that show the object. The features should increase significantly during the interjection period (frames 12-15) for the four-frame interjection and object removal datasets.
IV-C1 Position 1 - Image Embeddings
@@
In all three feature plots, the baseline clean DAVIS dataset is relatively constant, as expected. The metric applied to the four-frame interjection dataset (red line) increases significantly during the interjection period (frames 12-15), as each interjection frame is compared to the prefix frames (frames 0-11). On average, each interjection frame is further from the last known frame containing the object than two frames from the same video are from each other.
The metrics applied to the context removal dataset (green line) follows an interesting pattern in the image embedding space. Ideally, these plots will be constant, like the clean DAVIS plots. Because the context is removed during the interjection period, the frames are very distant from earlier frames. The large jump in feature values in frame 12 reflect this frame dissimilarity. However, the object is still present. As such, the subsequent frames are being compared to the removed-context frames, resulting in a lower distance for the remainder of the interjection period. The short-term distance and ratio plots show an additional jump in frame 16 when the context returns.
The metrics applied to the object removal dataset (orange line) experiences the inverse effect. These plots should match the four-frame interjection plots with a large jump during the interjection period. Because the background context of the frames in the interjection period remain, the frame is generally similar during this period. This is shown by the lack of a large increase during the interjection period. Instead. the feature values increase slightly, representing the small change in frame caused by the removed object.
IV-C2 Position 4 - Object Pointer
@@
The object pointer space is the ideal representation of the object separated from its surroundings. Analysis of the object pointer can create a nearly complete picture of the segmented object. Further analysis of the observed features and additional plots for all positions are shown in Appendix B
V Conclusion
SAM 2, the current state-of-the-art video object segmentation model, employs a unique architecture that solely relies on cross-attention with memory. In order to improve real-world applicability of tracking, localization, and segmentation models, particularly on complex scenes, it is necessary to understand how each stage of the SAM 2 architecture interprets video. In this paper, Five new datasets with complex transforms were proposed, each testing the robustness of SAM 2 on different challenges. Five different observational positions at each major stage within the model were analyzed, and it was determined at which stage the model is able to discern the object of interest from obscuring objects and the surrounding context. Primarily, the object pointer provides unique insight into where the model is spatially focused when writing to memory and when subsequent frames perform memory cross-attention. By visualizing how each architectural structure impacts the overall video understanding, VOS development can improve real-world applicability despite complex cluttered scenes and video transformations.
Appendix A Analytical Methods
@@
In real-world applications, it is useful to distinguish frames containing an object from frames in which the object is not present [8]. The computed metric can be used to create a classifier to separate visible frames from artificially interjected frames without the object. In this paper, three features are explored that result from these computed regularized distances that may be used as a classifier. The nature of SAM 2 can be understood by quantifying the differences between clean and transformed videos in the equivalent observational positions.
A-A Zeroth-Order L2 Comparison
@@
A-B First-Order L2 Comparison
In certain cases, the most dissimilar frames from the same video may be less similar than two frames from different videos. As such, zeroth-order classification is imperfect. First-order classification looks at the ratio of the current zeroth-order distance to the previous distance. This allows the model to determine if an interjected frame is less similar to the original video than the original video is to itself, thus improving the classification accuracy. In this paper, the short-term L2 ratio as a feature of each stage is explored.
Appendix B Extended Results
B-A Position 1 - Image Embeddings
@@
B-B Position 2 - Pixel/Memory Features
@@
B-C Position 3 - Pixel/Prompt Embeddings
@@
B-D Position 4 - Object Pointer
@@
B-E Position 5 - Memory Features